Use a change of variable to find solutions for  \frac{\mathrm{d} y}{\mathrm{d} x}=sin^2(x-y).

Hint: Let u = x – y
Answer #1

oby Sun 2c7-y) - 0 Substitute Uzx-y (7-) 3862- hele la care este Now Put the value of u and dye in ② oyna Sin? (2-y) l-du 2 S20 tan (x-4)=xte fulx-y] 50 tan (x-4)= x+c) or is the reqp. Solution (2-y)? tom(X+6). - 9 = x - tan-(4+0)]